# Integrating Conductor Health into Dynamic Line Rating and Unit Commitment under Uncertainty

### v2

@@ II-A Heat balance equation
where $q_{c}$, $q_{r}$, and $q_{s}$ denote convective and radiative heat losses and solar heat gain, respectively, and $I^{2}R(\tau_{\text{avg}})$ is the Joule heating term. Among these, wind speed has the strongest influence on the DLR, followed by ambient temperature and solar radiation [7].
II-B Linear temperature approximation
The current–temperature relationship is convex under realistic operating conditions, as confirmed through parametric sweeps of an ACSR Finch conductor (current: 500–3500 A, wind speed: 0–8 m/s, ambient temperature: 0–45 °C). Concave regions appear only for extreme overheating ($\tau\geq 350$ °C, wind speed $\leq 2$ m/s), which are beyond the safe operating range for ACSR. Thus, a linear approximation provides a conservative and computationally efficient proxy. Figure 2 shows the linear fit of the HBE-derived temperature curve. Within the emergency loading range (100–150% of rated current [27]), the linear proxy yields a mean absolute error of 0.7% and a maximum error of 2.2%. Because the function is convex, the linearization is constructed to slightly overestimate temperature, ensuring conservativeness.
@@ II-B Linear temperature approximation
where $p_{l,t}$ is the line power flow. Because voltage is assumed constant in DC power flow, $A_{l,t}$ is scaled by $\frac{1}{\sqrt{3}V^{\text{Nom}}_{l}}$ from the current–temperature fit. The two inequalities account for bidirectional flow, ensuring $\tau_{l,t}$ is an upper bound on the actual conductor temperature.
Figure 1: Illustration of a linear fit to the current-conductor temperature relationship. The temperature from the IEEE HBE and the linear fit assuming ACSR Finch, a perpendicular wind speed of 2.7 m/s and ambient temperature of 9.4°C is shown. Conductor temperature limit is set as 95 °C; in practice, it varies depending on the transmission system owner [28].
Figure 2: An illustrative example of LoTS with a conductor with exposed temperatures: 105°C, 109°C, 112°C each for two hours. The path of total degradation calculation is shadowed on trajectories of different temperatures [19].
II-C Loss of tensile strength (LoTS) for ACSR lines
Long-term ETO causes conductor degradation, measured as the loss of tensile strength (LoTS). For aluminum strands, LoTS (%) is empirically modeled as [18]:
$L_{\mathrm{Al}}=100-A\,t^{-\frac{25.4}{d}(0.001\tau-0.095)},$ (3)
where $A=\min(100,\,134-0.24\tau)$, $\tau$ is conductor temperature (°C), $t$ is exposure time (h), and $d$ is conductor diameter (mm). For aluminum conductor steel-reinforced (ACSR) lines, the corresponding LoTS is approximated as [19]:
$L_{\text{ACSR}}=0.735\,L_{\text{Al}}.$ (4)
II-D Piecewise depreciation cost for overhead lines
Operation of ACSR lines beyond 95°C accelerates tensile strength loss; thus, LoTS can be translated into economic depreciation. The ratio of the accumulated LoTS to an end-of-life value (typically 10% [29, 17]) represents the remaining conductor value. The cost of one hour ETO at a given temperature is calculated by incorporating the capital cost of a transmission line into a linear depreciation model. When the LoTS limit is reached, the line is assumed to lose its value. The depreciation is calculated as follows:
$C^{\text{Dep}}_{l,t}(\tau)=B_{l}\cdot S_{l}\cdot D_{l}\cdot\frac{L_{\text{ACSR},t}(\tau)-L_{\text{ACSR},t-1}}{10},$ (5)
where $B_{l}$ is the initial cost factor of line $l$ (\$/MVA-km), $S_{l}$ its capacity (MVA), and $D_{l}$ its length (km). Subscript $t$ denotes the time of interest. This formulation implies that $C^{\text{Dep}}_{t}(\tau)$ varies with the state of the conductor as $L_{\text{ACSR},t(\cdot)}$ is state-dependent. While the depreciation model in (5) can be applied in post-hoc evaluation, this research aims to integrate it into the UC problem with DLR which involves the real-time decision-making. Thus, we approximate the expected cost of a one-hour ETO by formulating the 1-hour ETO costs at different temperatures as a piecewise function prior to solving the daily CHA-UC problem. This function is convex, as higher temperature yield steeper depreciation slopes, and is calculated separately for each conductor. Figure 3 depicts the depreciation cost function and its piecewise counterpart for an old transmission line with many ETO histories (old) and a transmission line with less ETO histories (new).
Figure 3: Depreciation cost piecewise functions for new and old ACSR Finch. The piecewise functions lay above the exact depreciation cost.
III Conductor health-aware unit commitment
@@
III-A Baseline UC model
Before presenting the CHA-UC framework, the deterministic baseline model for DLR and SLR is introduced. Day-ahead UC and real-time operation is reflected in the model. The UC problem fixes commitment decisions day before the actual operation, using day-ahead forecasts and considering physical constraints of generators. The enforcement of UC constraints lead to a more comprehensive benchmark of results [14]. The operation step adjusts the generation of available generators under ramping constraints and fixed commitments to account forecast error.
@@ III-A Baseline UC model
which is the sum of production cost, startup cost and the shut down cost. $\Xi:=\{\mathbf{g,p,u,}\,\bm{\theta,\delta}\}$
The day-ahead UC constraints, hereafter referred to as the UC constraints, are adapted from the formulation of Morales-España et al. [30]:
@@
where $\Xi:=\{\mathbf{g}^{\mathrm{adj}},\mathbf{p}^{\mathrm{adj}},\mathbf{r}^{\mathrm{+,-}},\bm{\theta}^{\mathrm{adj}}\}$. Reserve activations are penalized through $C_{i}^{\mathrm{r}\pm}$, and the quadratic term captures nonlinear production costs. This objective represents the operator’s total adjustment cost in real time. The constraints of the operation stage are as follows:
@@
III-B Conductor health-aware unit commitment
Given the linear fit and the piecewise depreciation cost function in Section II, the objective function of the CHA-UC day-ahead problem is as follows:
@@ III-B Conductor health-aware unit commitment
The objective includes the operation cost from (6) and the scenario stochastic costs: the expected conductor depreciation cost, expected load shed costs and the expected reserve activation costs. The decision variables are identical to the baseline. The additional term in (11) represents the expected values from considering stochasticity. The sample average approximation was used to calculate the expected value. $C^{\text{Dep}}_{l,t}$ is a piecewise function of the depreciation cost of ETO in Section II-D, implemented using SOS2 constraints.
@@
$\begin{split}&\sum_{l\mid s(l)=b}\hskip-4.2679pt\hat{p}^{\text{adj}}_{l,t,s}-\sum_{l\mid r(l)=b}\hskip-4.2679pt\hat{p}^{\text{adj}}_{l,t,s}+\sum_{i\in\mathcal{I}_{b}}\hat{g}^{\text{adj}}_{i,t,s}+\hskip-4.2679pt\sum_{w\in\mathcal{W}_{b}}\hskip-4.2679pt(\hat{P}^{\rm wind}_{w,t}\\ &-\hat{p}^{\text{cur,adj}}_{w,t}-\omega_{w,t,s})=P^{\text{d}}_{b,t}-\hat{p}^{\text{shed,adj}}_{b,t,s},\quad\forall b\in\mathcal{B},t\in\mathcal{T},\end{split}$ (13f)
@@
The additional conductor temperature constraints enable the conductor temperature to be reflected in the objective function.
III-C CHA-UC Timeline
@@
The CHA-UC timeline is shown in Fig. 4. a UC problem is solved 18:00 the day before, and an operation problem is solved real-time (or hour-ahead). The loading-temperature linear fit should be updated everyday before the UC problem is solved to reflect the most recent conductor condition based on the latest forecasts. After operation, the depreciation cost function is updated for the next optimization period.
IV Case Study
Two case studies are performed on the synthetic Texas 123-bus backbone transmission system (TX-123BT) [31] using 2021-2022 weather data. Case Study I runs an annual simulation with uncertain DLR forecasts (with errors) and perfect wind power (WP) forecasts (no errors) to illustrate CHA-UC. Case Study II evaluates CHA-UC in a more realistic setting with uncertainty in both WP and DLR forecasts to analyze their correlation effects.
TABLE I: Summarized results of Case Study I and Case Study II. The best result for each cost is in bold.
@@ IV Case Study
CHA 14.32 0.13 0.27 11.5 14.68 14.70 1.88 1.62 0.53 104.6 18.75
Spring DLR 14.17 1.38 10.3 15.55 13.73 2.71 4.05 8.56 79.6 29.05
SLR 14.57 0.26 83.8 14.89 14.15 2.28 2.91 5.52 94.9 24.87
CHA 29.96 0.13 0.26 20.1 30.39 29.75 2.91 2.31 0.31 49.6 35.36
Summer DLR 29.59 1.26 3.4 30.89 28.43 4.12 5.26 5.78 31.9 43.63
SLR 29.92 0.85 64.5 30.80 28.79 3.67 4.20 5.65 33.8 42.36
CHA 22.10 0.08 0.04 14.6 22.26 22.00 2.55 1.36 0.16 74.6 26.13
Fall DLR 21.82 0.47 18.2 22.32 20.87 3.66 3.48 5.58 47.6 33.60
SLR 22.23 0.06 108.2 22.33 21.13 3.27 2.52 5.30 81.3 32.38
CHA 18.53 0.06 0.23 4.1 18.84 18.65 2.07 0.90 0.10 63.7 21.75
Winter DLR 18.39 0.94 4.4 19.33 17.71 3.01 3.01 6.33 41.7 30.08
SLR 18.86 51.2 18.89 18.14 2.61 2.10 4.42 60.0 27.28
CHA 21.25 0.09 0.20 12.7 21.57 21.30 2.36 1.55 0.28 73.2 25.55
Annual DLR 21.02 1.01 9.1 22.05 20.21 3.38 3.96 6.56 50.3 34.14
SLR 21.42 0.30 77.1 21.74 20.61 2.96 2.94 5.23 67.5 31.77
IV-A Preliminaries
IV-A1 Simulation Settings
@@
In all case studies, a set of 21 lines was selected for DLR application [^1] Lines were screened by solving an annual DCOPF with SLR and selecting those with the highest dual values of the line-flow constraint; lines longer than 100 km were excluded because their limits are dominated by stability.. All other lines used SLR limits. Each 345 kV line was modeled as ACSR Finch, and ampacity was adjusted to fit the TX-123BT system through a line multiplier. All optimization models were implemented in Julia/JuMP [32] and solved with Gurobi. The LoTS limit was set to 10% (Section II-D). The lifetime cost factor $B_{l}$ was based on a 3190 MVA, 400 kV AC line with a cost of 750 GBP/MVA-km [33]. After excluding tower and loss components, the adjusted cost factor ($B_{l}$) was \$936/MVA-km. The loading-temperature linear fit was sampled from 95 °C and 195 °C, and the conductor depreciation piecewise function was sampled every 10 °C from 95 °C. For DLR security margin, line limits were constrained as $K_{l,t}\hat{\overline{F}}^{\mathrm{DLR}}_{l,t}=\max(0.8\hat{\overline{F}}^{\mathrm{DLR}}_{l,t},\,\underline{F}^{\mathrm{SLR}}_{l})$. The load profile of 2021 given in the TX-123BT test system was used, assuming no uncertainties in the load forecast. The value of lost load, $C^{\text{VoLL}}$, is assumed to be \$3500/MWh following MISO before 2025/05 [34]. Reserve activation costs followed [35]: $C_{i}^{\mathrm{r}+}$ equals three times, and $C_{i}^{\mathrm{r}-}$ equals 0.5 times, the linear cost coefficient of generator $i$.
IV-A2 Data Collection
@@
Weather data for 2021–2022 were obtained from the NOAA’s high resolution rapid refresh model [36], which provides 3 km spatial resolution and 1 h temporal resolution. Forecasts issued at 18:00 the previous day were used as day-ahead inputs, while analysis data represented actual operation. Weather data from 2022 were used to generate stochastic scenarios, and 2021 data were used for annual simulations. WP profiles were derived from wind speeds using the standard power curve [37], adjusted to hub height via roughness-length correction. Uniform turbine specifications (cut-in and cut-out speeds) were assumed. For each transmission line, the hourly minimum DLR across all 3 km $\times$ 3 km grid cells intersecting the line corridor was used as the representative value. The parameters in (1) followed IEEE [26], with solar heat gain fixed at the worst-case condition for Texas (latitude 31°, 1000 W/m2, day 180, and maximum projected area). The conductor thermal limit was set to 95°C.
((a))
((b))
@@
Figure 6: (Case Study I) Annual scatter plot of temperatures of line 192 (Line with biggest temperature difference) obtained through post-hoc evaluation. 20-day moving average (MA) lines are drawn above. Meteorogical seasons are colored in the background. A zoom-in of thermal overload temperatures during periods when the MA of SLR is higher than CHA-UC, is provided.
IV-A3 Scenario Generation for CHA-UC
@@
A total of 20 scenarios were generated from the historical data. Figure 5 shows that the resulting distribution slightly overestimates the empirical tails, ensuring adequate tail-risk representation. The correlation structure was also well preserved—the average empirical correlation was 0.04, compared to 0.05 across the generated scenarios—demonstrating effective reproduction of inter-asset dependencies.
IV-B Case study I - Value of Conductor Health Awareness

### v4

@@ 2.4 Heat balance equation
where $q_{c}$, $q_{r}$, and $q_{s}$ denote convective and radiative heat losses and solar heat gain, respectively. $I^{2}R(\tau_{\text{avg}})$ is the Joule heating term. Among these, wind speed has the strongest influence on the DLR, followed by ambient temperature and solar radiation Wallnerstrom2015. Additional information on the details of the HBE can be found in IEEE738.
2.5 Linear temperature approximation
[…]
@@ 2.5 Linear temperature approximation
where $p_{l,t}$ is the line power flow. Since voltage is assumed constant in the DC power flow, $A_{l,t}$ is scaled by $\frac{1}{\sqrt{3}V^{\text{Nom}}_{l}}$ from the current–temperature fit. The two inequalities account for bidirectional flow, ensuring $\tau_{l,t}$ is an upper bound on the actual conductor temperature. The use of a linear approximation makes it unnecessary to explicitly express the absolute value of current in optimization, simplifying the conductor temperature model.
3 Conductor health-aware unit commitment
@@
3.1 Baseline UC model
Before presenting the CHA-UC framework, the deterministic baseline model for DLR and SLR is introduced. Day-ahead UC and real-time operation is reflected in the model. The UC problem fixes commitment decisions day before the actual operation, using day-ahead forecasts and considering physical constraints of generators. The enforcement of UC constraints lead to a more comprehensive benchmark of results Park2018. The operation step adjusts the generation of available generators under ramping constraints and fixed commitments to account forecast error.
@@ 3.1 Baseline UC model
which is the sum of production cost, startup cost and the shut down cost. $\Xi:=\{\mathbf{g,p,u,}\,\boldsymbol{\theta,\delta}\}$
The day-ahead UC constraints, hereafter referred to as the UC constraints, are adapted from the formulation of Morales-España et al. Morales2013:
@@
where $\Xi:=\{\mathbf{g}^{\mathrm{adj}},\mathbf{p}^{\mathrm{adj}},\mathbf{r}^{\mathrm{+,-}},\boldsymbol{\theta}^{\mathrm{adj}}\}$. Reserve activations are penalized through $C_{i}^{\mathrm{r}\pm}$, and the quadratic term captures nonlinear production costs. This objective represents the operator’s total adjustment cost in real time. The constraints of the operation stage are as follows:
@@
3.2 Conductor health-aware unit commitment
Given the linear fit and the piecewise depreciation cost function in Section 2, the objective function of the CHA-UC day-ahead problem is as follows:
@@ 3.2 Conductor health-aware unit commitment
The objective includes the operation cost from (9) and the scenario stochastic costs: the expected conductor depreciation cost, expected load shed costs and the expected reserve activation costs. The operation cost and the expected conductor depreciation cost use same units (\$), since the conductor depreciation model is designed to be incorporated to the UC problem. This eliminates the need for any additional scaling or conversion coefficient, which simplifies the optimization and improves interpretability. The decision variables are identical to the baseline. The additional term in (14) represents the expected values from considering stochasticity. The sample average approximation was used to calculate the expected value. $C^{\text{Dep}}_{l,t}$ is a piecewise function of the depreciation cost of ETO in Section 2.3, implemented using SOS2 constraints.
@@
$\begin{split}&\sum_{l\mid s(l)=b}\hskip-4.2679pt\hat{p}^{\text{adj}}_{l,t,s}-\sum_{l\mid r(l)=b}\hskip-4.2679pt\hat{p}^{\text{adj}}_{l,t,s}+\sum_{i\in\mathcal{I}_{b}}\hat{g}^{\text{adj}}_{i,t,s}+\hskip-4.2679pt\sum_{w\in\mathcal{W}_{b}}\hskip-4.2679pt(\hat{P}^{\rm wind}_{w,t}\\ &-\hat{p}^{\text{cur,adj}}_{w,t}-\omega_{w,t,s})=P^{\text{d}}_{b,t}-\hat{p}^{\text{shed,adj}}_{b,t,s},\quad\forall b\in\mathcal{B},t\in\mathcal{T},\end{split}\hskip-8.53581pt$ (16f)
@@
The additional conductor temperature constraints enable the conductor temperature to be reflected in the objective function. Analogous to equations (16h)-(16i), equations (19a)-(19b) enforce $\tau_{l,t}$ to reflect the conductor temperature before the operation stage.
3.3 CHA-UC Timeline
@@
The CHA-UC timeline is shown in Figure 4. A UC problem is solved 18:00 the day before, and an operation problem is solved real-time (or hour-ahead). The loading-temperature linear fit should be updated everyday before the UC problem is solved to reflect the most recent conductor condition based on the latest forecasts. After operation, the depreciation cost function is updated for the next optimization period.
4 Case Study
Two case studies are performed on the synthetic Texas 123-bus backbone transmission system (TX-123BT) Lu2025 using 2021-2022 weather data. Case Study I runs an annual simulation with uncertain DLR forecasts (with errors) and perfect wind power (WP) forecasts (no errors) to illustrate CHA-UC. Case Study II evaluates CHA-UC in a more realistic setting with uncertainty in both WP and DLR forecasts to analyze their correlation effects.
Table 1: Summarized results of Case Study I and Case Study II. The best result for each cost is in bold.
@@ 4 Case Study
Spring CHA 14.29 0.09 0.23 11.7 14.60 15.29 1.72 0.77 0.49 116.8 18.26
DLR 14.17 1.31 10.3 15.47 13.73 2.71 3.98 11.67 77.2 32.09
SLR 14.57 0.12 108.4 14.69 14.15 2.29 2.85 8.88 92.8 28.17
QRF 14.25 0.69 29.9 14.94 13.90 2.55 3.41 8.32 82.5 28.17
Summer CHA 29.99 0.12 0.13 11.9 30.21 30.79 2.56 1.39 0.14 52.7 34.87
DLR 29.63 0.95 3.4 30.57 28.48 4.12 5.15 10.81 28.9 48.56
SLR 29.96 0.46 64.8 30.42 28.84 3.68 4.28 8.35 38.9 45.14
QRF 30.35 0.19 32.94 30.55 29.63 3.77 4.61 6.32 34.9 44.33
Fall CHA 22.08 0.07 0.01 24.8 22.15 22.83 2.27 0.97 0.01 69.2 26.09
DLR 21.85 0.49 18.3 22.34 20.89 3.65 3.43 12.68 47.7 40.65
SLR 22.26 0.04 108.2 22.30 21.28 3.26 2.78 7.27 77.5 34.60
QRF 22.01 0.07 44.7 22.08 21.16 3.48 3.06 7.02 51.9 34.71
Winter CHA 18.50 0.05 0.13 1.1 18.67 19.28 1.91 0.61 0.00 75.6 21.81
DLR 18.39 0.65 4.4 19.04 17.72 3.01 3.02 8.70 48.6 32.44
SLR 18.86 0.00 50.9 18.86 18.15 2.61 2.06 4.98 57.0 27.80
QRF 18.44 0.27 4.3 18.71 17.81 2.94 2.81 6.01 40.8 29.56
Annual CHA 21.23 0.08 0.12 12.5 21.44 22.07 2.12 0.94 0.16 78.6 25.29
DLR 21.04 0.85 9.1 21.89 20.23 3.38 3.90 10.98 50.6 38.49
SLR 21.44 0.16 83.4 21.60 20.63 2.96 3.00 7.39 66.6 33.99
QRF 21.29 0.31 28.2 21.60 20.65 3.19 3.48 6.92 52.7 34.24
4.1 Preliminaries
4.1.1 Simulation Settings
@@
In all case studies, a set of 21 lines was selected for DLR application [^1] Lines were screened by solving an annual DCOPF with SLR and selecting those with the highest dual values of the line-flow constraint; lines longer than 100 km were excluded because their limits are dominated by stability.. All other lines used SLR limits. Each 345 kV line was modeled as ACSR Finch, and ampacity was adjusted to fit the TX-123BT system through a line multiplier. All optimization models were implemented in Julia/JuMP Lubin2023 and solved with Gurobi. The LoTS limit was set to 10% (Section 2.3). The lifetime cost factor $B_{l}$ was based on a 3190 MVA, 400 kV AC line with a cost of 750 GBP/MVA-km Parsons2012. After excluding tower and loss components, the adjusted cost factor ($B_{l}$) was \$936/MVA-km. The PDF modeling line failure $f(\cdot)$ was modeled as a normal distribution following $\mu=0.1,\sigma=0.05$. The loading-temperature linear fit was sampled from 95 °C and 150 °C, and the conductor depreciation piecewise function was sampled every 10 °C from 95 °C. Conductor health was initialized to fit a mean LoTS rate of 1% (a sensible rate from the results of Adomah2000) by creating synthetic historical ETO exposures from a normal distribution. The ETO exposure data is passed on between different horizons to update the conductor state properly. For DLR security margin, line limits were constrained as $K_{l,t}\hat{\overline{F}}^{\mathrm{DLR}}_{l,t}=\max(0.8\hat{\overline{F}}^{\mathrm{DLR}}_{l,t},\,\underline{F}^{\mathrm{SLR}}_{l})$. The load profile of 2021 given in the TX-123BT test system was used, assuming no uncertainties in the load forecast. The value of lost load, $C^{\text{VoLL}}$, is assumed to be \$3500/MWh following MISO before 2025/05 Kirschen2018. Reserve activation costs followed Dvorkin2025: $C_{i}^{\mathrm{r}+}$ equals three times, and $C_{i}^{\mathrm{r}-}$ equals 0.5 times, the linear cost coefficient of generator $i$.
4.1.2 Data Collection
@@
Weather data for 2021–2022 were obtained from the NOAA’s high resolution rapid refresh model Benjamin2016, which provides 3 km spatial resolution and 1 h temporal resolution. Forecasts issued at 18:00 the previous day were used as day-ahead inputs, while analysis data represented actual operation. Weather data from 2022 were used to generate stochastic scenarios, and 2021 data were used for annual simulations. WP profiles were derived from wind speeds using the standard power curve Giorsetto1983, adjusted to hub height via roughness-length correction. Uniform turbine specifications (cut-in and cut-out speeds) were assumed. For each transmission line, the hourly minimum DLR across all 3 km $\times$ 3 km grid cells intersecting the line corridor was used as the representative value. The parameters in (7) followed IEEE IEEE738, with solar heat gain fixed at the worst-case condition for Texas (latitude 31°, 1000 W/m2, day 180, and maximum projected area). The conductor thermal limit was set to 95°C.
@@
4.1.3 Scenario Generation for CHA-UC
@@
A total of 20 scenarios were generated from the historical data. Figure 5 shows that the resulting distribution slightly overestimates the empirical tails, ensuring adequate tail-risk representation. The correlation structure was also well preserved–the average empirical correlation was 0.04, compared to 0.05 across the generated scenarios–demonstrating effective reproduction of inter-asset dependencies.
Figure 6: (Case Study I) Annual scatter plot of temperatures of line 192 (Line with biggest temperature difference) obtained through post-hoc evaluation. 20-day moving average (MA) lines are drawn above. Meteorogical seasons are colored in the background. (a) The temperatures of DLR, SLR and CHA-UC are compared. A zoom-in of thermal overload temperatures during periods when the MA of SLR is higher than CHA-UC, is provided. (b) The temperatures of QRF and CHA-UC are compared.
4.1.4 Benchmark Model
A quantile regression forest (QRF) model was used to benchmark the performance of the CHA-UC model. QRF is an ensemble method that generates multiple decision trees from the training data and produces quantile forecasts from the forest output. The model setting was adopted from Dupin et al. Dupin2019. The forest was created using 1-year data from 2022. The confidence level was set to 90%, and the overestimation frequency was 7.4%. The QRF forecasts are conservative and capture the overall temporal pattern of the actual DLR. The quantile trajectories move with the observations, although sharp peaks are smoothed as a result of the ensemble-based quantile estimation. For the QRF-UC simulation, the DLR line constraints (11h) were replaced with QRF forecasts.
4.2 Case study I - Value of Conductor Health Awareness